OpenStudy (anonymous):

a diner serves lunch for $6. an average of 90 lunches are sold per week. a market survey indicates that for each 25-cent increase in the price, the number of customers per week will decrease by 3. a) determine a function R(x) that models the weekly revenue after x 25-cent increases in the price. b) what price should the diner charge to maximize the weekly revenue? what is the maximum revenue?

OpenStudy (anonymous):

please help me!

OpenStudy (jamesj):

So R(0) = # of lunches sold after NO prices increases. What is that equal to?

OpenStudy (anonymous):

6

OpenStudy (jamesj):

No, how many lunches are sold today.

OpenStudy (anonymous):

wait i mean 90

OpenStudy (jamesj):

Right. R(1) = # lunches sold after 1 25 cent increase. What's that equal to?

OpenStudy (anonymous):

not sure sorry

OpenStudy (jamesj):

The question says: "a market survey indicates that for each 25-cent increase in the price, the number of customers per week will decrease by 3." Hence if the price increases by 25 cents, how many customers will there be?

OpenStudy (jamesj):

There is 90. If you take 90 and decrease it by 3, what do you get?

OpenStudy (anonymous):

87

OpenStudy (jamesj):

Yes, so R(1) = 90 - 3 Now what if there are two 25 cent increases? What is R(2) equal to?

OpenStudy (anonymous):

84

OpenStudy (jamesj):

Right, R(2) = 90 - 2*3 Hence in general, if there are x 25 cent increases, how many customers will there be R(x) ?

OpenStudy (jamesj):

R(0) = 90 R(1) = 90 - 1*3 R(2) = 90 - 2*3 ...

OpenStudy (anonymous):

R(x) = 90-3x ??

OpenStudy (jamesj):

correct.

OpenStudy (anonymous):

so how about part b?

OpenStudy (jamesj):

Now you want to maximize revenue. Revenue = (Price) x (# customers) Write Q(x) for revenue, P(x) for price and R(x) for customers. You just found an expression for R(x). Find now an expression for P(x). Then Q(x) = P(x).R(x)

OpenStudy (jamesj):

Then you want to use calculus to maximize the function Q(x).

OpenStudy (anonymous):

can you tell me what the answers are and then i can ask you if i dont understand it?

OpenStudy (jamesj):

Nope. Use the same procedure we just used to find R(x) to find P(x). P(0) = $6.00 if we have one price increase the price is what? That's P(1). Keep on going until you figure out what P(x) is.

OpenStudy (anonymous):

i dont get it

OpenStudy (jamesj):

Where are you lost? Do you understand where R(x) came from?

OpenStudy (jamesj):

R(x) = 90 - 3x

OpenStudy (anonymous):

right i get that

OpenStudy (jamesj):

Do you agree that revenue = price * units sold ?

OpenStudy (anonymous):

yes

OpenStudy (jamesj):

Hence if we want to maximize Revenue, we need to find an expression for Revenue. We only have one piece of the equation: units sold. Now we need the other piece of the equation price. Write P(x) as the function for price. What is P(0) = price with no price increases?

OpenStudy (anonymous):

6

OpenStudy (jamesj):

And what is P(1) ?

OpenStudy (anonymous):

6.25

OpenStudy (jamesj):

Hence what is P(x) in general?

OpenStudy (anonymous):

6+.25x ?

OpenStudy (jamesj):

correct, P(x) = 6 + 0.25x. hence what is the equation for revenue as a function of x. Write that as Q(x). Q(x) = what?

OpenStudy (anonymous):

not sure sorry!

OpenStudy (jamesj):

Revenue = Price * Units sold.

OpenStudy (anonymous):

90 (6+.25x)

OpenStudy (jamesj):

Price = P(x) units sold = R(x) Hence, revenue = Q(x) = ... what?

OpenStudy (anonymous):

90-3x (6+.25x) ?

OpenStudy (jamesj):

Q(x) = R(x).P(x) = (90 - 3x)(6 + 0.25x)

OpenStudy (jamesj):

Now use calculus to find the maximum for Q(x). I.e., you're solving the problem: what is the number of prices increases x such that revenue Q(x) is maximized.

OpenStudy (anonymous):

how do i find that....?

OpenStudy (jamesj):

Do you know calculus?

OpenStudy (anonymous):

im in precalc and my teacher didnt teach us this. can you just please tell me?

OpenStudy (jamesj):

In which case, your teacher is expecting you to use other things you know about quadratic equations. Q(x) = R(x).P(x) = (90 - 3x)(6 + 0.25x) Expand that expression out and put it in standard form for a quadratic: ax^2 + bx + c Then use what you know about quadratic equations to find its maximum. In this case, the maximum will be at the vertex of the graph of the quadratic.

OpenStudy (anonymous):

the question wants price and the maxium revenue...when i find the vertex on the graph is that for price or maximum revenue

OpenStudy (jamesj):

for maximum revenue because the function you are graphing is Q(x), the expression for revenue.

OpenStudy (anonymous):

ok and what is maximum price then?

OpenStudy (jamesj):

That will be the value of x corresponding to where Q(x) is maximized.

OpenStudy (jamesj):

I mean it will be P(x) for the x where Q(x) is maximized.

OpenStudy (anonymous):

what does that mean

OpenStudy (jamesj):

x is the number of price increases. The price for x price increases is P(x) = 6 + 0.25x.

OpenStudy (jamesj):

You are maximizing revenue, Q(x), as a function of x, that is, as a function of the number of 25 cent price increases. That's the way the problem has been set up for you.

OpenStudy (anonymous):

i got .03 for maximum revenue? that doesnt make sense

OpenStudy (jamesj):

No. What is your expression for Q(x) written as a quadratic in standard form?

OpenStudy (jamesj):

Q(x) = R(x).P(x) = (90 - 3x)(6 + 0.25x)

OpenStudy (anonymous):

-.75x^2 + 4.5x +540

OpenStudy (jamesj):

exactly, good. Now, for what value of x is that maximized?

OpenStudy (anonymous):

now i got 2.9 ??

OpenStudy (jamesj):

for what value of x is ax^2 + bx + c have a vertex?

OpenStudy (jamesj):

*does

OpenStudy (anonymous):

i dont know!!!

OpenStudy (jamesj):

x = -b/2a

OpenStudy (anonymous):

which is 3 which is what i got (2.9)

OpenStudy (jamesj):

Hence Q(x) = -.75x^2 + 4.5x +540 has a vertex when \[ x = \frac{-4.5}{2(-1.5)} = 3 \]

OpenStudy (jamesj):

This is the value of x for which Q(x) is maximized.

OpenStudy (jamesj):

That corresponds to 3 prices increases of 25 cents.

OpenStudy (anonymous):

thats the rmaximum revenue?????

OpenStudy (jamesj):

The maximum revenue is the value of Q(x) when x = 3.

OpenStudy (anonymous):

546. 75

OpenStudy (anonymous):

and what is the price then

OpenStudy (jamesj):

Yes, Q(3) = $546.75. You tell me what the price is. What does x = 3 mean? We've written down what it means for price about four times already.

OpenStudy (anonymous):

6.75 ?

OpenStudy (jamesj):

yes. P(x) = 6 + 0.25x. Hence P(3) = $6.75

OpenStudy (anonymous):

thanks for working through this with me.

OpenStudy (jamesj):

Candy, Whenever I have a new hard problem like this one. After I've figured it out, I take a blank piece of paper and rewrite the solution. Then I do it again When I can do that from a blank piece of paper without looking at the earlier solutions, I know I understand it. This is a great way to reinforce what you've learnt and I recommend you try it. I know learning new things is painful. But it's very satisfying when you get it.

OpenStudy (anonymous):

thanks. is there any way to simple this down:\[x^3-5x^2-2x+24+5IX^2-5IX-30I\]

OpenStudy (anonymous):

i got that from multiplying (x^2 -x -6) (x-4+5i)